OpenStudy (anonymous):

Is point (1, 4) on the line represented by the equation x + 2y = 5?

OpenStudy (anonymous):

If f(x) = -x + 12, find f(2) and give the ordered pair.

OpenStudy (ajprincess):

for the first question plug in the values x=1 nd y=4 in x+2y. if u get 5 then the point lies on the line represented by the equation x+2y=5. Does that help? @Mariahs

OpenStudy (anonymous):

a little

OpenStudy (ajprincess):

hmm. when u plug in the values x=1 nd y=4 in x+2y u will get 1+2(4). 1+2(4)=?

OpenStudy (ajprincess):

@Mariahs

OpenStudy (anonymous):

huh

OpenStudy (ajprincess):

can u plz tell me what is 1+2(4)=?